d^2q/dt^2+16dq/dt=60
In an L-C circuit, L = 1 henry, C = 1
16 farad and E(t) = 60 volt.
Solce using laplace

dt dt represents the capacitor charge at anytime t.
Given q(0)= 0and q' ( 0)=i(0)=0 find:
7.1 the charge q on the capacitor at anytime t. (7)
7.2 the current i at anytime t.
1
Expert's answer
2021-07-18T16:09:03-0400
"s^2Y(s)+16sY(s)=\dfrac{60}{s}"

"Y(s)=\dfrac{60}{s^2(s+16)}"

"\dfrac{60}{s^2(s+16)}=\dfrac{A}{s+16}+\dfrac{B}{s}+\dfrac{C}{s^2}"

"=\dfrac{As^2+Bs(s+16)+C(s+16)}{s^2(s+16)}"

"s=0:C=\dfrac{15}{4}"

"s=-16:A=\dfrac{15}{64}"

"B=-A=-\dfrac{15}{64}"

7.1

"q(t)=\dfrac{15}{4}t-\dfrac{15}{64}+\dfrac{15e^{-16t}}{64}"

7.2

"i(t)=\dfrac{15}{4}-\dfrac{15e^{-16t}}{4}"